OpenStudy (anonymous):

What is the value of f(x) = 9x when x = -2?

Vocaloid (vocaloid):

substitute -2 for x f(x) = 9x f(-2) = 9(-2) = ?

OpenStudy (anonymous):

-18

OpenStudy (anonymous):

9(-2)=f(x) f(x)=-18

OpenStudy (anonymous):

derp

Vocaloid (vocaloid):

yes, -18 is the answer

OpenStudy (anonymous):

but the choices does not say -18

Vocaloid (vocaloid):

uh, are you sure you typed your function in correctly? did you mean f(x) = 9^x?

OpenStudy (anonymous):

(a)1/81 (b)81 (c)1/18 (d)18

OpenStudy (anonymous):

9^x

OpenStudy (anonymous):

\[1/81\]

OpenStudy (anonymous):

A then

Vocaloid (vocaloid):

yeah. f(-2) = 9^(-2) = 1/81

OpenStudy (anonymous):

thank you soo much

OpenStudy (anonymous):

cause -2 makes the 9^2 go upside down to make the -2 positive

OpenStudy (anonymous):

:3 np